# A mini-review on combinatorial solutions to the Marcus–Lushnikov irreversible aggregation

### v2

@@ I Introduction
The next successful attempt at analytical description of the irreversible aggregation was the stochastic approach by Marcus Marcus_1968 , later studied and developed by Lushnikov Lushnikov_1978 ; Lushnikov_2004 ; Lushnikov_2005 ; Lushnikov_2011 . It may be regarded as a microscopic model of the deterministic Smoluchowski model, both describing spatially averaged systems (spatially invariant) and being quantitatively similar for sufficiently large systems. However, the Marcus–Lushnikov approach gives a more detailed description of the process than Smoluchowski. It allows us to ask questions—such as the probability distribution of particles of given size at a given time or the variability between realizations—that are beyond the scope of the deterministic Smoluchowski rate equations vanDongen_1987_fluct ; vanDongen_1987_fluct_II ; Leyvraz_2022 . Another key difference concerns large clusters (large with respect to the system size). While the Smoluchowski equation allows for clusters of unbounded size and represents averages over all possible orders of aggregation events, the Marcus–Lushnikov model represents one possible sequence of coagulation events and restricts clusters to size less than a finite number (finite system size, NN). This makes it possible to study the distributions of clusters for any given size, including those similar to NN in both non-gelling and gelling systems Leyvraz_2022 . In general, this stochastic Marcus–Lushnikov model provides expected time-dependent size-distribution profiles (either by analytical solutions or averaging over many numerical simulations), which can be compared to solutions of the Smoluchowski model. It also gives information on the distribution of possible resulting states of the system, which has no equivalent in the Smoluchowski approach (cf. Sec. II.9).
@@ II.1 Initial assumptions
The system consists of NN monomers (the smallest units, monomer particles), some of which later join into clusters. $N=\text{const}$ is equivalent to the preservation of mass in the system. A single coagulation act occurs in one time step; therefore, the total number of clusters, $k$, at time step $t$ is
@@
Using the above relation, it can be shown that the distribution of the cluster numbers $z=n_{s}/N$ as a function of $t$ tends towards $)\delta(z-c_{s}(\tilde{t}$)). As a result, the only difference between the two frameworks is that the Smoluchowski theory provides a less detailed description than the Marcus–Lushnikov theory, as the former does not describe the full distribution of cluster numbers, and always involves a limit $N\to\infty$. In this way, for large values of NN, the distribution of $n_{s}$ is sharply peaked around the value given by the Smoluchowski theory. However, for small NN, for which the distribution does not reduce to a delta function, Marcus–Lushnikov theory provides a fuller view. What also follows is that the variance of the distribution is beyond the reach of the Smoluchowski theory, though it vanishes for the limit of $N\to\infty$ Leyvraz_2022 .
@@
However, at the cost of a somewhat complex relationship with real time, we obtain a different benefit. Suppose we have information on the initial number of clusters and a “snapshot” of the cluster size distribution at a particular point in time (thus, NN and $k$). In that case, the translation to the language of discrete time steps is straightforward because a particular number of clusters in the system clearly defines the time step in the combinatorial formalism (t=Nkt=N-k). Such an empirical procedure of comparing combinatorial predictions to the experimental data from a small discrete system has been shown in detail in Ref. Lepek_2021 .
@@ II.3 Probability distribution over state space
Figure 4: a) A graphical representation of the cluster selection process in the case of coagulation, in which the probability of two particles merging does not depend on their sizes. Our task is to create a particle of size $g$, i.e., consisting of $g$ monomers. The monomers are distinguishable (numbered for clarity). In the first step, we combine two out of $g$ available particles (here $g=4$). In the next step, we combine two further freely selected clusters. We merge the particles until we obtain a single cluster. In this way, we implement one of $x_{4}=4!3!/2^{3}=18$ ways to create a cluster of size 44. b) Diagram illustrating all possible growth histories of such a particle of size 44. The number of different growth histories, $x_{g}$, equals the number of different paths drawn by arrows and leading through different states of the diagram. c) A tree corresponding to the bold path in the diagram in b). Single particle growth history may be illustrated as a rooted tree, with leaf nodes standing for monomeric units, internal nodes representing the history-dependent transition states of the cluster, and the root node being the final step in the growth process. Each vertical line represents a state of the system at a particular time step. Illustration by the authors.
@@ II.4 Distributing monomers into clusters
In this section, we will calculate the number of ways to distribute monomers into the given number of clusters (i.e., partitioning the system of NN monomers into $k$ subsets) Fronczak_2018 . As both $s$ and $g$ denote cluster size, at this point, we will use $g$, which will prove convenient in subsequent derivations, cf. Eqs. (22)-(26). Let NN distinguishable objects be divided into $k$ non-empty and disjoint subsets of $c_{i}>0$ elements each, where $\sum^{k}_{i=1}{c_{i}=N}$. As we can choose $c_{1}$ in $\binom{N}{c_{1}}$ ways, $c_{2}$ in $\binom{N-c_{1}}{c_{2}}$ ways, there are
@@
For example, let us consider a system of $N=5$ monomers grouped into $k=2$ clusters of sizes 2 and 3. Remember that all monomers are distinguishable (i.e., we can assign numbers from 11 to $N=5$ to these particles). Possible grouping methods are presented in Fig. 3a. There are ten ways of doing so.
@@
To sum up this part, Eq. (11) gives the number of possible ways of partitioning NN monomers into $k$ clusters for a given state $\mathrm{\Omega}$ of the system. However, it implicitly assumes that all the clusters in the system arose at once, which is not valid, as only one merging act is allowed per time step. Coagulation acts corresponding to different clusters may alternate, increasing the number of ways in which a given (final) state can be created. This issue leads to the need to calculate the number of possible distributions of merging acts in time (see Sec. II.6).
@@ II.6 Distributing joining acts in time
$\begin{split}\left[\binom{t}{1}\binom{t-1}{1}\dots\binom{t-n_{2}+1}{1}\right]\dots\left[\binom{t-n_{2}}{2}\binom{t-n_{2}-2}{2}\dots\binom{t-n_{2}-2(n_{3}-1)}{2}\right]\dots{\\ }=\frac{t!}{{\left(1!\right)}^{n_{2}}{\left(2!\right)}^{n_{3}}\dots\left(\left(g-1\right)!\right)^{n_{g}}\dots}=t!\prod^{N-k+1}_{g=1}{\frac{1}{{\left(\left(g-1\right)!\right)}^{n_{g}}}}{=W_{t}(\mathrm{\Omega})}.\end{split}$ (14)
@@ II.7 Expressions for average number of clusters of given size and its standard deviation
It should also be noted that $\omega_{s}$ is a single value depending on the cluster size $s$ while $\omega_{g}$ is a sequence which is independent of $s$, where $g$ varies from 11 to Nk+1N-k+1 (i.e., to $t+1$). Clearly, for cluster sizes $s\geq t+2$ we have $\left\langle n_{s}\right\rangle=0$ as clusters of such sizes had no chance to arise.
@@ II.9 Probability distribution for number of clusters of given size
Up to this moment, we have been focused on finding the average number $\langle n_{s}\rangle$ of clusters of a given size $s$, Eq. (27), and its corresponding standard deviation $\sigma_{s}$, Eq. (29). In fact, these two quantities constitute the first and the second moment of the time-dependent probability distribution for the number of clusters of a given size, $P(n_{s},t)$. It is the probability that there are exactly $n_{s}$ clusters of size $s$ in the system consisting of NN monomers, in which there are k=Ntk=N-t clusters in total.
@@ III.1 Constant kernel, $K=\mathrm{const}$
Using the above result, we can use NsN-s instead of NN and k1k-1 instead of $k$ to obtain
@@
A comparison of the statistics obtained by the above equations to the results of numerical simulations of a system of $N=100$ monomers aggregating at a constant rate is presented in Fig. 7. What deserves emphasis, variance increases at long times as the set of possible system states widens. Later, when the system approaches its final state (t=N1t=N-1) with only a single cluster, variance quickly decreases as the number of possible states decreases. Also, observing values of $\sigma_{s}^{2}/\langle n_{s}\rangle$ close to 11 tells us that the actual fluctuations are of the same intensity as the mean; thus, we can expect that for particular sizes $g$, we may not find such clusters in the system at all.
@@ III.4 Linear–chain kernel, $K\propto(s^{-1}+g^{-1})^{\alpha}$
where, of course, $g$ varies from 11 to Nk+1N-k+1.
@@ III.5 Condensation kernel, $K\propto(A+s)(A+g)$
where $A\geq 0$ stands for an arbitrary process-dependent parameter. This kernel may be regarded as a generalization of the product kernel with AA being a factor that relatively decreases the rate of the process for larger particles, thus preventing the creation of a giant cluster (gel).
@@ III.7 Other kernels solved
Figure 11: Some kernels solved in the combinatorial approach may be transformed with the parameter AA to the form corresponding to the constant kernel. These are: linear-chain, $K\propto\left(s^{-1}+g^{-1}\right)^{\alpha}$, condensation, $K\propto(A+s)(A+g)$, and linear combination of constant and additive kernels, $K\propto A+s+g$. In the case of the latter two, by taking $A=0$, we obtain basic kernels, product and additive, respectively. In this way, explicit solutions to all “intermediate” processes between the product, additive, linear-chain, and the constant kernel were obtained. Of course, the above picture is only a schematic representation: exact dynamics throughout the aggregation process depend on the current values of AA, $s$, $g$, and the system size. For instance, for $K\propto A+s+g$, if AA is large, there is an early stage when the system is dominated by clusters of size $s\ll A$, and the dynamics are dominated by $K\propto A$. Later, when the average cluster size is $s\simeq A$, all terms in $K\propto A+s+g$ are equally significant. At later times, when the average cluster size is $s\gg A$, the dynamics are dominated by the term $K\propto s+g$, which may be observed for sufficiently large NN. For the linear-chain kernel, the branch $\alpha\to\infty$” symbolically represents processes with $\alpha>1$, for which increasing $\alpha$ lowers the relative probability of merging for large clusters. Precision of the theoretical predictions for the above-mentioned kernels varies from exact to approximate, see Sec. V.1 for details.
@@ IV.1 Using recursive expression for $x_{g}$
Using the recursive relation, Eq. (31), requires a starting value, $x_{1}$, which may be interpreted as the number of possible internal states of a cluster of size 11. As for all of the previous kernels, we will use $x_{1}=1$, which is a reasonable guess and has worked well up to this point. A given kernel $K$ used in Eq. (31) defines further numbers in the sequence, $x_{2}$, $x_{3}$, and so on.
@@ IV.2 Example 1: Protoplanetary aggregation
The formation of planets is a complex process that still needs to be better understood. In general, planet formation may be regarded as consisting of several stages. First, micrometer-sized cosmic dust grains aggregate into pebbles (mm/cm), then the creation of planetesimals (10–100 km) occurs. Planetesimals are solid objects thought to exist in protoplanetary disks. They can efficiently accrete leftover pebbles and directly form the cores of planets, i.e., protoplanets of sizes greater than 1000 km, which may later evolve into terrestrial planets or gas giants Levison_2015 ; Liu_2020 . The current approach to studying cosmic matter aggregation uses NN-body simulations as in Ref. Levison_2015 .
@@ IV.3 Example 2: Growth of aerosols
On the other hand, for large KnKn, the motion of particles becomes ballistic, and a limit for this regime may be found as
@@
and AA, $B$ and $C$ are empirically-determined parameters Rogak_1992 .
@@ 11 Exact Fronczak_2018 ; Leyvraz_2022
Table 3: Kernels solved in the combinatorial approach, i.e., where an explicit expression for $x_{g}$ was found. In the second column, we summarize the precision of the theoretical predictions given by the expression for $\langle n_{s}\rangle$, Eq. (27). The above summary is based on comparisons to the numerical simulations and on theoretical asymptotic analysis performed in Ref. Leyvraz_2022 . In the general case, the combinatorial approach yields approximate solutions that, when analyzed in the large system size limit, are incompatible with known results from the scaling theory (see the main text and Appendix C for details). However, the three basic kernels and their combinations (), due to their special scaling properties, behave differently from the general case. For the constant and additive kernels, it has been shown that the combinatorial solutions yield exact results both in the large-scale limit of $\langle n_{s}\rangle$ as well as for the full probability distribution as explicitly evaluated for a small system. Although not explicitly enumerated, the same shall apply for the combination $K=A+s+g$ Leyvraz_2022 . For the product kernel, the precision of the results drops after the gelling point but preserves the existence of the giant cluster. Asymptotic analysis has shown that the combinatorial solution for this case may be regarded as exact in the large-scale limit for the time before the gelling point, while it yields a known (inaccurate) solution after the gelling point. $(^{\star\star})$ In case of the linear-chain (LC) kernel, decreasing $\alpha$ means that the kernel form tends to the one of the constant kernel; thus, behavior and precision also tend to the behavior and precision of the constant kernel. The discrepancies for small-cluster sizes (with $\alpha=1$) were studied in Ref. Leyvraz_2022 as a result of the incompatibility with the scaling theory. $(^{\star\star\star})$ Analogously, for the condensation kernel, increasing AA means that the contribution of $s$ and $g$ decreases and the kernel behavior becomes more similar to the behavior of the constant kernel, resulting in increasing precision (of course, exact dynamics throughout the aggregation process depends on the current values of AA, $s$, $g$, and the system size). For $A=0$, the solution is identical to the one for $K=sg$, and all the concerns related to the product kernel apply.

### v4

@@ I Introduction
The next successful attempt at analytical description of the irreversible aggregation was the stochastic approach by Marcus Marcus_1968 , later studied and developed by Lushnikov Lushnikov_1978 ; Lushnikov_2004 ; Lushnikov_2005 ; Lushnikov_2011 . It may be regarded as a microscopic model of the deterministic Smoluchowski model, both describing spatially averaged systems (spatially invariant) and being quantitatively similar for sufficiently large systems. However, the Marcus–Lushnikov approach gives a more detailed description of the process than Smoluchowski. It allows us to ask questions—such as the probability distribution of particles of given size at a given time or the variability between realizations—that are beyond the scope of the deterministic Smoluchowski rate equations vanDongen_1987_fluct ; vanDongen_1987_fluct_II ; Leyvraz_2022 . Another key difference concerns large clusters (large with respect to the system size). While the Smoluchowski equation allows for clusters of unbounded size and represents averages over all possible orders of aggregation events, the Marcus–Lushnikov model represents one possible sequence of coagulation events and restricts clusters to size less than a finite number (finite system size, $N$). This makes it possible to study the distributions of clusters for any given size, including those similar to $N$ in both non-gelling and gelling systems Leyvraz_2022 . In general, this stochastic Marcus–Lushnikov model provides expected time-dependent size-distribution profiles (either by analytical solutions or averaging over many numerical simulations), which can be compared to solutions of the Smoluchowski model. It also gives information on the distribution of possible resulting states of the system, which has no equivalent in the Smoluchowski approach (cf. Sec. II.9).
@@ II.1 Initial assumptions
The system consists of $N$ monomers (the smallest units, monomer particles), some of which later join into clusters. $N=\text{const}$ is equivalent to the preservation of mass in the system. A single coagulation act occurs in one time step; therefore, the total number of clusters, $k$, at time step $t$ is
@@
Using the above relation, it can be shown that the distribution of the cluster numbers $z=n_{s}/N$ as a function of $t$ tends towards $)\delta(z-c_{s}(\tilde{t}$)). As a result, the only difference between the two frameworks is that the Smoluchowski theory provides a less detailed description than the Marcus–Lushnikov theory, as the former does not describe the full distribution of cluster numbers, and always involves a limit $N\to\infty$. In this way, for large values of $N$, the distribution of $n_{s}$ is sharply peaked around the value given by the Smoluchowski theory. However, for small $N$, for which the distribution does not reduce to a delta function, Marcus–Lushnikov theory provides a fuller view. What also follows is that the variance of the distribution is beyond the reach of the Smoluchowski theory, though it vanishes for the limit of $N\to\infty$ Leyvraz_2022 .
@@
However, at the cost of a somewhat complex relationship with real time, we obtain a different benefit. Suppose we have information on the initial number of clusters and a “snapshot” of the cluster size distribution at a particular point in time (thus, $N$ and $k$). In that case, the translation to the language of discrete time steps is straightforward because a particular number of clusters in the system clearly defines the time step in the combinatorial formalism (t=Nkt=N-k). Such an empirical procedure of comparing combinatorial predictions to the experimental data from a small discrete system has been shown in detail in Ref. Lepek_2021 .
@@ II.3 Probability distribution over state space
Figure 4: a) A graphical representation of the cluster selection process in the case of coagulation, in which the probability of two particles merging does not depend on their sizes. Our task is to create a particle of size $g$, i.e., consisting of $g$ monomers. The monomers are distinguishable (numbered for clarity). In the first step, we combine two out of $g$ available particles (here $g=4$). In the next step, we combine two further freely selected clusters. We merge the particles until we obtain a single cluster. In this way, we implement one of $x_{4}=4!3!/2^{3}=18$ ways to create a cluster of size $4$. b) Diagram illustrating all possible growth histories of such a particle of size $4$. The number of different growth histories, $x_{g}$, equals the number of different paths drawn by arrows and leading through different states of the diagram. c) A tree corresponding to the bold path in the diagram in b). Single particle growth history may be illustrated as a rooted tree, with leaf nodes standing for monomeric units, internal nodes representing the history-dependent transition states of the cluster, and the root node being the final step in the growth process. Each vertical line represents a state of the system at a particular time step. Illustration by the authors.
@@ II.4 Distributing monomers into clusters
In this section, we will calculate the number of ways to distribute monomers into the given number of clusters (i.e., partitioning the system of $N$ monomers into $k$ subsets) Fronczak_2018 . As both $s$ and $g$ denote cluster size, at this point, we will use $g$, which will prove convenient in subsequent derivations, cf. Eqs. (22)-(26). Let $N$ distinguishable objects be divided into $k$ non-empty and disjoint subsets of $c_{i}>0$ elements each, where $\sum^{k}_{i=1}{c_{i}=N}$. As we can choose $c_{1}$ in $\binom{N}{c_{1}}$ ways, $c_{2}$ in $\binom{N-c_{1}}{c_{2}}$ ways, there are
@@
For example, let us consider a system of $N=5$ monomers grouped into $k=2$ clusters of sizes 2 and 3. Remember that all monomers are distinguishable (i.e., we can assign numbers from $1$ to $N=5$ to these particles). Possible grouping methods are presented in Fig. 3a. There are ten ways of doing so.
@@
To sum up this part, Eq. (11) gives the number of possible ways of partitioning $N$ monomers into $k$ clusters for a given state $\mathrm{\Omega}$ of the system. However, it implicitly assumes that all the clusters in the system arose at once, which is not valid, as only one merging act is allowed per time step. Coagulation acts corresponding to different clusters may alternate, increasing the number of ways in which a given (final) state can be created. This issue leads to the need to calculate the number of possible distributions of merging acts in time (see Sec. II.6).
@@ II.6 Distributing joining acts in time
$\begin{split}\left[\binom{t}{1}\binom{t-1}{1}\dots\binom{t-n_{2}+1}{1}\right]\left[\binom{t-n_{2}}{2}\binom{t-n_{2}-2}{2}\dots\binom{t-n_{2}-2(n_{3}-1)}{2}\right]\dots{\\ }=\frac{t!}{{\left(1!\right)}^{n_{2}}{\left(2!\right)}^{n_{3}}\dots\left(\left(g-1\right)!\right)^{n_{g}}\dots}=t!\prod^{N-k+1}_{g=1}{\frac{1}{{\left(\left(g-1\right)!\right)}^{n_{g}}}}{=W_{t}(\mathrm{\Omega})}.\end{split}$ (14)
@@ II.7 Expressions for average number of clusters of given size and its standard deviation
It should also be noted that $\omega_{s}$ is a single value depending on the cluster size $s$ while $\omega_{g}$ is a sequence which is independent of $s$, where $g$ varies from $1$ to Nk+1N-k+1 (i.e., to $t+1$). Clearly, for cluster sizes $s\geq t+2$ we have $\left\langle n_{s}\right\rangle=0$ as clusters of such sizes had no chance to arise.
@@ II.9 Probability distribution for number of clusters of given size
Up to this moment, we have been focused on finding the average number $\langle n_{s}\rangle$ of clusters of a given size $s$, Eq. (27), and its corresponding standard deviation $\sigma_{s}$, Eq. (29). In fact, these two quantities constitute the first and the second moment of the time-dependent probability distribution for the number of clusters of a given size, $P(n_{s},t)$. It is the probability that there are exactly $n_{s}$ clusters of size $s$ in the system consisting of $N$ monomers, in which there are k=Ntk=N-t clusters in total.
@@ III.1 Constant kernel, $K=\mathrm{const}$
Using the above result, we can use NsN-s instead of $N$ and k1k-1 instead of $k$ to obtain
@@
A comparison of the statistics obtained by the above equations to the results of numerical simulations of a system of $N=100$ monomers aggregating at a constant rate is presented in Fig. 7. What deserves emphasis, variance increases at long times as the set of possible system states widens. Later, when the system approaches its final state (t=N1t=N-1) with only a single cluster, variance quickly decreases as the number of possible states decreases. Also, observing values of $\sigma_{s}^{2}/\langle n_{s}\rangle$ close to $1$ tells us that the actual fluctuations are of the same intensity as the mean; thus, we can expect that for particular sizes $g$, we may not find such clusters in the system at all.
@@ III.4 Linear–chain kernel, $K\propto(s^{-1}+g^{-1})^{\alpha}$
where, of course, $g$ varies from $1$ to Nk+1N-k+1.
@@ III.5 Condensation kernel, $K\propto(A+s)(A+g)$
where $A\geq 0$ stands for an arbitrary process-dependent parameter. This kernel may be regarded as a generalization of the product kernel with $A$ being a factor that relatively decreases the rate of the process for larger particles, thus preventing the creation of a giant cluster (gel).
@@ III.7 Other kernels solved
Figure 11: Some kernels solved in the combinatorial approach may be transformed with the parameter AA to the form corresponding to the constant kernel. These are: linear-chain, $K\propto\left(s^{-1}+g^{-1}\right)^{\alpha}$, condensation, $K\propto(A+s)(A+g)$, and linear combination of constant and additive kernels, $K\propto A+s+g$. In the case of the latter two, by taking $A=0$, we obtain basic kernels, product and additive, respectively. In this way, explicit solutions to all “intermediate” processes between the product, additive, linear-chain, and the constant kernel were obtained. Of course, the above picture is only a schematic representation: exact dynamics throughout the aggregation process depend on the current values of $A$, $s$, $g$, and the system size. For instance, for $K\propto A+s+g$, if AA is large, there is an early stage when the system is dominated by clusters of size $s\ll A$, and the dynamics are dominated by $K\propto A$. Later, when the average cluster size is $s\simeq A$, all terms in $K\propto A+s+g$ are equally significant. At later times, when the average cluster size is $s\gg A$, the dynamics are dominated by the term $K\propto s+g$, which may be observed for sufficiently large NN. For the linear-chain kernel, the branch $\alpha\to\infty$” symbolically represents processes with $\alpha>1$, for which increasing $\alpha$ lowers the relative probability of merging for large clusters. Precision of the theoretical predictions for the above-mentioned kernels varies from exact to approximate, see Sec. V.1 for details.
@@ IV.1 Using recursive expression for $x_{g}$
Using the recursive relation, Eq. (31), requires a starting value, $x_{1}$, which may be interpreted as the number of possible internal states of a cluster of size $1$. As for all of the previous kernels, we will use $x_{1}=1$, which is a reasonable guess and has worked well up to this point. A given kernel $K$ used in Eq. (31) defines further numbers in the sequence, $x_{2}$, $x_{3}$, and so on.
@@ IV.2 Example 1: Protoplanetary aggregation
The formation of planets is a complex process that still needs to be better understood. In general, planet formation may be regarded as consisting of several stages. First, micrometer-sized cosmic dust grains aggregate into pebbles (mm/cm), then the creation of planetesimals (10–100 km) occurs. Planetesimals are solid objects thought to exist in protoplanetary disks. They can efficiently accrete leftover pebbles and directly form the cores of planets, i.e., protoplanets of sizes greater than 1000 km, which may later evolve into terrestrial planets or gas giants Levison_2015 ; Liu_2020 . The current approach to studying cosmic matter aggregation uses $N$-body simulations as in Ref. Levison_2015 .
@@ IV.3 Example 2: Growth of aerosols
On the other hand, for large $Kn$, the motion of particles becomes ballistic, and a limit for this regime may be found as
@@
and $A$, $B$ and $C$ are empirically-determined parameters Rogak_1992 .
@@ 11 Exact Fronczak_2018 ; Leyvraz_2022
Table 3: Kernels solved in the combinatorial approach, i.e., where an explicit expression for $x_{g}$ was found. In the second column, we summarize the precision of the theoretical predictions given by the expression for $\langle n_{s}\rangle$, Eq. (27). The above summary is based on comparisons to the numerical simulations and on theoretical asymptotic analysis performed in Ref. Leyvraz_2022 . In the general case, the combinatorial approach yields approximate solutions that, when analyzed in the large system size limit, are incompatible with known results from the scaling theory (see the main text and Appendix C for details). However, the three basic kernels and their combinations (), due to their special scaling properties, behave differently from the general case. For the constant and additive kernels, it has been shown that the combinatorial solutions yield exact results both in the large-scale limit of $\langle n_{s}\rangle$ as well as for the full probability distribution as explicitly evaluated for a small system. Although not explicitly enumerated, the same shall apply for the combination $K=A+s+g$ Leyvraz_2022 . For the product kernel, the precision of the results drops after the gelling point but preserves the existence of the giant cluster. Asymptotic analysis has shown that the combinatorial solution for this case may be regarded as exact in the large-scale limit for the time before the gelling point, while it yields a known (inaccurate) solution after the gelling point. $(^{\star\star})$ In case of the linear-chain (LC) kernel, decreasing $\alpha$ means that the kernel form tends to the one of the constant kernel; thus, behavior and precision also tend to the behavior and precision of the constant kernel. The discrepancies for small-cluster sizes (with $\alpha=1$) were studied in Ref. Leyvraz_2022 as a result of the incompatibility with the scaling theory. $(^{\star\star\star})$ Analogously, for the condensation kernel, increasing AA means that the contribution of $s$ and $g$ decreases and the kernel behavior becomes more similar to the behavior of the constant kernel, resulting in increasing precision (of course, exact dynamics throughout the aggregation process depends on the current values of $A$, $s$, $g$, and the system size). For $A=0$, the solution is identical to the one for $K=sg$, and all the concerns related to the product kernel apply.